OpenStudy (anonymous):

x/90=7/9... how can i solve for x

OpenStudy (anonymous):

Just cross-multiply 9*x = 90*7 9x = 630 x = 630/9 = ?

OpenStudy (anonymous):

x=70?

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

LOL! Thanks nin :)

OpenStudy (anonymous):

just type it on here?

OpenStudy (anonymous):

sorry im new to this...

OpenStudy (anonymous):

thanks alot @Fall12-13 !!!